### v1

# Random Dehn function of groups

Abstract.
In this note, we study the notion of random Dehn function and compute an asymptotic upper bound for finitely presented acylindrically hyperbolic groups whose Dehn function is at most polynomial with integer exponent. By showing that in these cases the random Dehn function is strictly smaller than the usual Dehn function we confirm Gromov’s intuition albeit in a different model.
Key words and phrases:
Dehn functions, filling invariants, random walks, Markov chains, acylindrically hyperbolic
2020 Mathematics Subject Classification:
Primary:20P05, 20F65, 20F69. Secondary:20F67
1. Introduction
Gromov’s seminal program of classifying groups up to quasi-isometries has been a central problem in geometric group theory. Among the quasi-isometry invariants that have emerged, the Dehn function of a finitely presented group has played a central role. The Dehn function of a finitely generated group $G$ is a complexity measure of a direct attack of the word problem in $G$: it provides an upper bound on the area of a word of length at most $n$, that is, on the number of relations that must be applied to a word in the generators of $G$, in order to prove that it represents the trivial element in $G$; the bound is expressed as a function of the length of the word. Geometrically, it provides an upper bound on the area of filling discs for a loop of length at most $n$ in a simply connected Riemannian manifold on which $G$ acts geometrically; the bound is given in terms of the length of the loop. An outstanding feature of Dehn functions is that they characterize hyperbolic groups as precisely those finitely presented groups having subquadratic Dehn function [Gro92, Ols91, Bow95].
The Dehn function of a finitely presented group, by definition, captures the worst-case complexity of the word problem in the group. However, for many algorithmic problems it is natural to inquire about the average-case complexity. Gromov defined [Gro92, 5.A6.c] the averaged Dehn function of a group $G$ as the average of the area over all words of length at most $n$ that represents the identity in $G$, and claimed that in several cases the asymptotic growth of the averaged Dehn function should be strictly smaller than that of the standard Dehn function. Young [You08] confirmed Gromov’s claim for the family of nilpotent groups.
Using another model of filling function, which we call random Dehn function, see Definition 2.10, we confirm Gromov’s intuition by showing that for finitely presented acylindrically hyperbolic groups whose usual Dehn function is at most polynomial, the random Dehn function grows strictly slower than the usual Dehn function. Roughly speaking, the random Dehn function of a group $G$ is defined as the expected values of areas in $G$. Intuitively, this is done by considering a random walk $(Z_{n})_{n}$ on $G$, which we think as a path on $G$. We then join the extreme points of this path, using a (quasi-geodesic) combing $\alpha$ for $G$, to obtain a closed loop on $G$. The random Dehn function $R\delta_{(G,(Z_{n}),\alpha)}$ of $G$ is then defined by taking the expected values of areas of loops constructed in this way.
Our main results, confirming Gromov’s intuition, are the following, with the usual order on functions which we recall later.
Theorem A.
Let $G$ be a finitely presented acylindrically hyperbolic group with Dehn function $\delta_{G}$ at most polynomial. Let $(Z_{n})_{n}$ be a simple random walk on $G$. For every quasi-geodesic combing $\alpha$ of $G$, we have that
$R\delta_{(G,(Z_{n}),\alpha)}(n)\preccurlyeq\frac{n}{\log(n)}\delta_{G}(\log(n)).$
In particular, $R\delta_{(G,Z_{n},\alpha)}(n)\preccurlyeq\delta_{G}(n)$.
We note that such a model of random Dehn function has been considered by Sisto in [Sis17, Theorem 4.10] in the case of relatively hyperbolic groups. Sisto shows that the random Dehn function of a hyperbolic group, relative to proper subgroups with at most polynomial Dehn funtion, is $\preccurlyeq n\delta(\log(n))$ where $\delta$ is the maximum of the Dehn functions of the peripheral subgroups.
Random walks belong to the larger class of stochastic processes called tame quasi-homogeneous Markov chains, whose definition we recall in Section 2. These process are invariant under quasi-isometries. In the case of tame Markov chains, we obtain the following result.
Theorem B.
Let $G$ be one of the following:
  • A non-elementary hyperbolic group;
  • a non-elementary relatively hyperbolic group;
  • an acylindrically hyperbolic $3$-manifold group;
  • the fundamental group of a graph of groups with nilpotent undistorted edge groups and at most polynomial Dehn function
Let $(w_{n}^{\ast})$ be a tame Markov chain on $G$ and $\alpha$ any quasi-geodesic combing of $G$, then we have that
$R\delta_{(G,(w^{*}_{n}),\alpha)}(n)\preccurlyeq\frac{n}{\log n}\delta_{G}(\log(n)).$
In particular, $R\delta_{(G,(w^{p}_{n}),\alpha)}(n)\prec\delta_{G}(n)$.
A subfamily of tame Markov chains, which is still invariant by quasi-isometry and includes simple random walks are tame quasi-homogeneous Markov chains. If we consider these, then we obtain the following result.
Theorem C.
Let $G$ be one of the following:
  • the mapping class group of a finite type surface;
  • an extra-large Artin group;
  • or more generally a well-behaved hierarchically hyperbolic group (in the sense of [GHP+23]).
Let $(w_{n}^{\ast})$ be a tame quasi-homogeneous Markov chain on $G$. Then, for any quasi-geodesic combing $\alpha$ of $G$ we have that:
$R\delta_{(G,(w^{*}_{n}),\alpha)}(n)\preccurlyeq n\log n.$
In particular, $R\delta_{(G,(w^{p}_{n}),\alpha)}(n)\prec\delta_{G}(n)$.
The proof of both of these theorems relies on deviation inequalities for the appropriate stochastic processes. Roughly, these inequalities state that the probability that the distance between the position of a random walk (or tame quasi-homogeneous Markov chain) after $k$ steps and a quasi-geodesic between the starting point and the position after $n$ steps is at least $t$ decays exponentially in $t$, see Theorem 3.2 for the exact statement. We then use these inequalities to deduce that our sample path coming from the stochastic process, with high probability, stays close to a quasi-geodesic which is enough to bound the area of the filling and hence give an estimate for the random Dehn function.
It is worth noting that another quasi-isometry invariant based on tame Markov chains has been considered by the second named author and A. Sisto in [GS23, GSon]. Instead of looking at the Dehn function of generic points, there the authors look at the divergence of generic points.
Acknowledgments
The authors thank the organisers of the thematic program on geometric group theory held at the Montreal Centre de Recherches Scientifiques in Spring 2023, where part of this work took place. The authors also thank Kunal Chawla and Alessandro Sisto for interesting discussions. The first author gratefully acknowledge previous funding by the DFG 281869850 (RTG 2229). The second author was supported by the EPSRC-UKRI studentship EP/V520044/1 and the LMS Early Career Fellowship.
2. Preliminaries
In this section, we recall some notions from geometric group theory as well as from the theory of random walks and Markov chains. We then define the notion of random Dehn function of a group.
2.1. Acylindrically hyperbolic groups
We recall the definition of an acylindrically hyperbolic groups, which is a generalisation of the notion of a hyperbolic group and we refer to [Osi16] and the references therein for more properties and examples of this class of groups.
Definition 2.1.
We say that a group acts acylindrically on a metric space $X$ if for every $R\geq 0$, there exist $N>0$, $L>0$ such that for every $x,y\in X$ with $d_{X}(x,y)\geq L$, we have
|$|\{g\in G\ |\ d_{X}(x,gx)\leq R,\ d_{X}(y,gy)\leq R\}|\leq N.$
A group is acylindrically hyperbolic if it is non-virtually cyclic and acts acylindrically on a hyperbolic geodesic metric space with unbounded orbits.
Example 2.1.
([Osi16]) The following groups are acylindrically hyperbolic
  • Non-elementary hyperbolic groups;
  • Non virtually cyclic relatively hyperbolic groups with proper peripheral subgroups;
  • All but a finite number of mapping class groups of connected oriented surfaces;
  • $Out(F_{n})$ for $n\geq 2$;
  • Non virtually cyclic groups acting properly on a CAT(0) space with a rank-1 element.
2.2. Markov chains
We usually denote a Markov chain on a group by $(w_{n}^{\ast})_{n}$, where $w_{n}^{p}$ denotes the position of the Markov chain starting at $p$ after $n$ steps and the quantity $\mathbb{P}[w_{n}^{o}=g]$ denotes the probability of getting from $o$ to $g$ in $n$ steps. We recall some definitions from [GS22] and [GHP+23] and refer to these for more details.
Definition 2.2.
For a finitely generated group $G$ with generating set $S$, with word metric $d_{S}$, we say that a Markov chain $(w_{n}^{\ast})$ on $G$ is tame if it satisfies the following conditions:
  1. (1)
    Bounded jumps: There exists a $K>0$ such that for all $g\in G,n\in\mathbb{N}$ we have that $d_{S}(w_{n}^{p},w^{p}_{n+1})\leq K$,
  2. (2)
    Non-amenability: there exists $A>0$ and $\rho<1$ such that for all $g,h\in G$ and $n\geq 0$ we have $\mathbb{P}[w_{n}^{g}=h]\leq A\rho^{n}$,
  3. (3)
    Irreducibility: For each $s\in G$ there exists constants $\epsilon_{s},K_{s}>0$ such that for all $g\in G$ we have $\mathbb{P}[w_{k}^{g}=gs]\geq\varepsilon_{s}$ for some $k\leq K_{s}$.
Definition 2.3.
For a bijection $f:G\rightarrow H$ and a Markov chain $(w^{\ast}_{n})$ on $G$, there is a natural Markov chain in $H$ induced by $f$ called the push-forward Markov chain on $H$, which is denoted by $f_{\#}(w^{p}_{n})$. This is the Markov chain such that $\mathbb{P}[f_{\#}(w^{o}_{n})=h]=\mathbb{P}[w_{n}^{o}=f^{-1}(h)]$ for all $h\in H$, $o\in G$ and $n\geq 0$.
Another property of Markov chains that we consider is the following.
Definition 2.4.
A Markov chain $(w^{p}_{n})_{n}$ in $G$ is quasi-homogeneous if for some $D>0$ it has the following property. For every $g,h\in G$ there is a bijection $f:G\rightarrow G$ such that $f(g)=h$ and for all $n\in\mathbb{N}$ and all $q\in G$ we have that $f_{\#}(w^{q}_{n})=(w_{n}^{f(q)})$.
The following result shows that appropriate random walks, including simple random walks, are tame quasi-homogeneous Markov chains.
Lemma 2.5.
([GS22, Lemma 2.8] [GHP+23]) Let $G$ be a non-amenable group and let $(z_{n})_{n}$ be a random walk on $G$ driven by $\mu$ such that the support of $\mu$ is finite and generates $G$ as a semigroup. Then $(z_{n})_{n}$ is a tame quasi-homogeneous Markov chain.
2.3. Dehn and random Dehn functions
Definition 2.6.
Let $S$ non-empty set and $F(S)$ the free group on $S$. Let $G$ be a group and suppose it is given by a finite presentation $\mathcal{P}=\langle S|R\rangle$. We denote by $|\cdot|_{S}$ the word-length with respect to $S$ and by $d_{S}$ the induced metric. We say that a word $w$ in $S$ is null-homotopic in $G$ if it represents the identity in $G$ and write $w=_{\mathcal{P}}1$. We define the area of a null-homotopic word $w$ in $G$ with respect to $\mathcal{P}$ as the positive integer defined by
$\operatorname{Area}_{\mathcal{P}}(w):=\min\{\ell\in\mathbb{N}\mid w=_{\mathcal{P}}\prod_{i=1}^{\ell}a_{i}r_{i}^{\epsilon_{i}}a_{i}^{-1},\ \text{with}\ a_{i}\in F(S),\ r_{i}\in R,\ \text{and}\ \epsilon_{i}\in\{\pm 1\}\}.$
The Dehn function of $\mathcal{P}$ is the function $\delta_{\mathcal{P}}:\mathbb{N}\rightarrow\mathbb{N}$ defined as
$\delta_{\mathcal{P}}(n):=\max\{\operatorname{Area}_{\mathcal{P}}(w)\mid w=_{\mathcal{P}}1,\ |w|_{S}\leq n\}.$
Changing the presentation of $G$ does not alter the asymptotic behaviour of the Dehn function of the presentations. Concretely, the Dehn function of $G$ is well-defined up to $\asymp$- equivalence of functions  [^1] Recall, two functions $f,g:[0,\infty)\rightarrow[0,\infty)$ are said to be $\asymp$-equivalent if $f\preccurlyeq g$ and $g\preccurlyeq f$, where $f\preccurlyeq g$ means that there exists constants $A,B>0$ and $C,D,E\geq 0$ such that $f(n)\leq Ag(Bn+C)+Dn+E$ for all $n\geq 0$., so we write $\delta_{G}=\delta_{\mathcal{P}}$. One can show that the Dehn function of a groups is a well-defined quasi-isometry invariant up $\asymp$-equivalence (for details see [Alo90] and [BH99]).
For a group $G$ generated by $S$, we can assume without loss of generality that $S$ is symmetric, namely that $S=S^{-1}$. There is a natural surjective map, called evaluation map, from the free monoid generated by $S$, denoted by $S^{\ast}$, to $G$ defined by sending each element in $S^{\ast}$, a word in $S$, into the element it represents in $G$.
Definition 2.7.
A morphism of monoids $\alpha:G\rightarrow S^{\ast}$ is called a combing for $G$ if it is a (set-theoretic) right inverse of the evaluation map. Geometrically, given $g\in G$ we can identify $\alpha(g)$ with a continuous path $\alpha_{g}\colon[0,l]\rightarrow\operatorname{Cay}(G,S)$ connecting $1$ to $g$ which we call the combing line of $g$.
In other words, a combing of $G$ is a choice of normal form for elements in $G$.
Definition 2.8.
Let $D\geq 1$. We say that a combing $\alpha\colon G\rightarrow S^{\ast}$ is a $D$-quasi-geodesic combing if for every $g\in G$ the combing line $\alpha_{g}$ is a $(D,D)$-quasi-geodesic. We simply say that $\alpha$ is a quasi-geodesic combing if there is some $D\geq 1$ such that $\alpha$ is a $D$-quasi-geodesic combing.
Definition 2.9.
Let $G$ be a finitely presented groups with finite presentation $\mathcal{P}=\langle S|R\rangle$ and $\alpha:G\rightarrow S^{\ast}$ be a combing for $G$. For $x_{0},\ldots,x_{n}\in G$ let
$\alpha_{i}=\alpha(x_{i+1}x_{i}^{-1})\quad\text{for}\ 0\leq i\leq n-1,\ \text{and}\quad\alpha_{n}=\alpha(x_{n}).$
For the null-homotopic word in $S$ defined as $W_{x_{1},\ldots,x_{n}}=\alpha_{1}\cdot\ldots\cdot\alpha_{n-1}\cdot\alpha_{n}^{-1}$ we denote
$\text{Fill}_{\mathcal{P},\alpha}(x_{1},\ldots,x_{n})=\operatorname{Area}_{\mathcal{P}}(W_{x_{1},\ldots,x_{n}}).$
Definition 2.10 (Random Dehn function).
Let $G$ be a finitely presented group with finite presentation $\mathcal{P}=\langle S|R\rangle$ and let $\alpha:G\rightarrow S^{\ast}$ be a quasi-geodesic combing for $G$. Let be $(w_{n}^{\ast})$ a stochastic process on $G$. For every starting point $o\in G$ of the stochastic process set
$x_{o,1}=w_{0}^{o},\ldots,x_{o,n}=w_{n}^{o},\quad\alpha_{o,i}=\alpha(x_{o,i+1}x_{o,i}^{-1})\quad\text{for}\ 0\leq i\leq n-1,\ \text{and}\quad\alpha_{o,n}=\alpha(x_{o,n}).$
The random Dehn function of $\mathcal{P}$ with respect to $(w_{n}^{\ast})$ and $\alpha$ is the function $R\delta_{(\mathcal{P},(w_{n}^{\ast}),\alpha)}\colon\mathbb{N}\rightarrow\mathbb{N}$ defined as
$R\delta_{(\mathcal{P},(w_{n}^{\ast}),\alpha)}(n)=\sup_{o\in G}\mathbb{E}[\text{Fill}_{\mathcal{P},\alpha}(x_{o,1},\ldots,x_{o,n})].$
The stochastic process $(w^{\ast}_{n})_{n}$ that we consider is either a random walk driven by an admissible measure, or a tame Markov chain. The reason for looking at tame quasi-homogeneous Markov chains instead of just random walks is that together with quasi-geodesic combings these make the notion of random Dehn function invariant under quasi-isometries, as we now show.
Proposition 2.11.
Let $G$ be a non-amenable group and $H$ be a group quasi-isometric to $G$. Then for all tame Markov chains $(z_{n}^{*})_{n}$ on $G$ and quasi-geodesic combing $\alpha$ on $G$, there exists a tame Markov chain $(w_{n}^{*})_{n}$ on $H$ and quasi-geodesic combing $\beta$ on $H$ such that for all $p\in G$
$R\delta_{(G,(z_{n}^{\ast}),\alpha)}(n)\asymp R\delta_{(H,(w_{n}^{\ast}),\beta)}(n).$
Proof.
By [Why99, Theorem 4.1], any quasi-isometry between finitely generated non-amenable groups is at bounded distance from a bijective quasi-isometry $f\colon G\to H$. We let $w_{n}^{*}=f_{\#}(z_{n}^{*})$ be the push-forward of $z_{n}^{*}$ by $f$ as defined in 2.3, then by [GS22, Lemma 2.8] and [GHP+23, Remark 2.9] this new stochastic process is a tame quasi-homogeneous Markov chain on $H$. Let $\beta=f\circ\alpha$, i.e. for all $x,y\in H$ let $\beta(x,y)=f(\alpha(f^{-1}(x),f^{-1}(y))$. The rest of the proof follows directly from the fact that the usual Dehn function is a quasi-isometric invariant. ∎
3. Bounding the random Dehn function
We first start with a result that shows that if a Markov chain makes linear progress then sub-walks of length at least logarithmic length also make linear progress. This is a generalisation of a result from [Sis17].
Lemma 3.1.
(Generalisation of [Sis17, Lemma 4.5]) Let $(w^{*}_{n})$ be a Markov chain on $G$ that makes linear progress with exponential decay in $G$. Then for each $r\geq 1$, there exists $C_{3}$ so that for each $n\geq 1$ and $o\in G$ we have:
$\mathbb{P}\Big{[}\exists i,j\leq n:|i-j|\geq C_{3}\log(n),d_{X}(w^{o}_{i},w^{o}_{j})\leq|i-j|/C_{3}\Big{]}\leq C_{3}n^{-r}.$
Proof of Lemma 3.1.
The proof is very similar to the proof in the case of a simple random walk proved in [Sis17, Lemma 4.5]. Let $K_{1}$ be the constant from linear progress with exponential decay of the Markov chain, in $G$. That is to say, let $K_{1}$ be such that $\mathbb{P}\Big{[}d_{S}(g,w_{k}^{o})<k/K_{1}\Big{]}\leq K_{1}e^{-k/K_{1}}.$ Let $r\geq 1$ and let $C_{3}=(r+1)K_{1}$. We have, for all $o\in G$ and $i<j$:
$\mathbb{P}\Big{[}d_{X}(w^{o}_{i},w^{o}_{j})\leq|j-i|/K_{1}\Big{]}=\sum_{g\in G}\mathbb{P}\Big{[}d_{X}(g,w^{g}_{j-i})\leq|j-i|/K_{1}\Big{|}w_{i}^{p}=g\Big{]}\mathbb{P}\Big{[}w_{i}^{p}=g\Big{]}\leq K_{1}e^{-|j-i|/K_{1}},$
where we used the Markov property. Hence, by the choice of $C_{3}$ we get
$\begin{split}\mathbb{P}\Big{[}\exists i,j\leq n:|i-j|\geq C_{3}\log(n),d_{X}(w^{o}_{i},w^{o}_{j})\leq|i-j|/C_{3}\Big{]}&\leq\sum_{i=0}^{n}\left(\sum_{l\geq C_{3}\log(n)}K_{1}e^{-l/K_{1}}\right)\\ &\leq K_{1}n\cdot n^{-C_{3}/K_{1}}\sum_{t\geq 0}e^{-t/K_{1}}\\ &\leq C_{3}n^{-r}.\end{split}$
The following result is a direct consequence of the deviation inequalities from [MS20, Theorem 1.1] and [GS22, Theorem 1.4]
Theorem 3.2 (Consequence of [GS22, MS20]).
We consider one of the following set ups.
  • Let $(w^{*}_{n})_{n}$ be a simple random walk on an acylindrically hyperbolic group $G$,
  • Let $(w^{*}_{n})_{n}$ be a tame Markov chain on a group satisfying [GS22, Assumption 1 or Assumption 2], for example a hyperbolic group, or relatively hyperbolic group or the fundamental group of a graph of groups with nilpotent undistorted edge groups,
  • Let $(w^{*}_{n})_{n}$ be a tame quasi-homogenenous Markov chain on a well-behaved hierarchically hyperbolic group, for example the mapping class group of a finite-type surface or an extra-large Artin group.
Then for all $D>0$, there exists a constant $C_{1}$ such that the following holds for all $n\in\mathbb{N}$
$\mathbb{P}\left[\forall k\leq n\colon\sup_{\alpha\in QG_{(D,D)}(1,Z_{n})}d_{S}(Z_{k},\alpha)\geq l\right]\leq C_{1}ne^{-l/C_{1}}$
the supremum is taken over the set $QG_{(D,D)}(x,y)$ of all $(D,D)$-quasi-geodesics from $x$ to $y$ in $G$.
Proof.
This follows immediately from the fact that in each of these set ups we have deviation inequalities by [MS20, Theorem 1.1] and [GS22, Theorem 1.4] (where for the last bullet point this is true via [GHP+23, Theorem 3]), hence summing up with a union bound over all $k\leq n$ gives the desired result. ∎
Theorem 3.3.
Let $G$ be a finitely presented group with presentation $\mathcal{P}=\langle S|R\rangle$ and with Dehn function $\delta_{G}$ at most polynomial. Let $(w_{n}^{*})_{n}$ be a tame Markov chain on $G$ that satisfies the conclusion of Theorem 3.2. Then, for all quasi-geodesic combings $\alpha$ of $G$, we have that
$R\delta_{(G,(w_{n}^{*}),\alpha)}(n)\preccurlyeq\frac{n}{\log(n)}\delta_{G}(\log(n)).$
Proof of Theorem 3.3.
We show that for all $D>0$, there exists a constant $C>0$ such that for all $D$-quasi-geodesic combings $\alpha$ of $G$ we have that
$R\delta_{(G,(w_{n}^{*}),\alpha)}(n)\leq C\frac{n}{\log(n)}\delta_{G}(\log(n))+Cn+C.$
Let $d$ be the degree of a polynomial bounding (from above) the Dehn function of $G$, the existence of $d$ is guaranteed by our assumptions. Let $(w_{n}^{*})_{n}$ be one of the above stochastic processes on the corresponding group $G$, the important part is that these satisfy the conclusion of Theorem 3.2. Let $o\in G$, $n\in\mathbb{N}$, and $\alpha_{o,w^{0}_{n}}$ be the quasi-geodesic line from $o$ to $w^{o}_{n}$, which we assume to be a $D$-quasi-geodesic.
We now provide a bound which is uniform over $o\in G$. Let $C_{1}$ be the constant from Theorem 3.2 associated to $D$, which we increase to make sure that $n^{d}C_{1}n^{1-C_{1}}\to 0$. Let $\mathcal{A}$ be the event
``$\forall k\leq n:d_{G}(\alpha_{o,w^{0}_{n}},w_{k}^{o})\leq C_{1}^{2}\log(n)^{\prime\prime}.$
By Theorem 3.2 the event $\mathcal{A}$ holds with probability $\mathbb{P}\Big{[}\mathcal{A}\Big{]}\geq 1-C_{1}n^{1-C_{1}}\to 1$.
We now work under the assumption that the event $\mathcal{A}$ holds, i.e. the event $\forall k\leq n:d_{S}(\alpha_{o,w^{0}_{n}},w^{o}_{k})\leq C_{1}^{2}\log(n)$. Hence, for all $k\leq n$, we can choose discrete geodesics $\gamma_{k}$ from $w^{o}_{k}$ to $\alpha_{o,w^{0}_{n}}$ of length at most $C^{2}_{1}\log(n)$.
Let $p_{k}=\alpha_{o,w^{0}_{n}}\cap\gamma_{k}$. For two points $x,y\in\alpha_{o,w^{0}_{n}}$, we write $\ell(\alpha_{o,w^{0}_{n}}|_{[x,y]})$ to denote the length, along $\alpha_{o,w^{0}_{n}}$, between $x$ and $y$.
Since $G$ is non-amenable, we get that a tame Markov chain makes linear progress in $G$ (by the non-amenability condition of tameness) and we can therefore use Lemma 3.1 to get that subwalks of logarithmic length also make linear progress in the Cayley graph of $G$ with respect to the chosen generating set.
Let $C_{3}$ be as in Lemma 3.1 for $r=d+1$ ($d$ being the polynomial degree bounding the Dehn function $\delta_{G}$ of $G$) and for all $i\geq 0$, define integers $k_{i}=\lceil 100iC_{1}^{2}C_{3}D^{2}\log(n)\rceil$ and let $t$ be such that $k_{t}\leq n$ but $k_{t+1}>n$. Define $\mathcal{B}$ to be the complementary event to the defining event of the probability in Lemma 3.1, i.e. $\mathcal{B}$ is the event
[…]
By Lemma 3.1, we have that $\mathbb{P}\big{[}\mathcal{B}^{C}\big{]}\leq C_{3}n^{-k}$ and so we are able to disregard the event $\mathcal{B}^{C}$. Hence, we now assume that $\mathcal{B}$ holds.
Claim 1.
If the events $\mathcal{A}$ and $\mathcal{B}$ both hold, then for all $k_{i+1}<n$ we have $\ell(\alpha_{o,w^{0}_{n}}|_{[o,p_{k_{i+1}}]})-\ell(\alpha_{o,w^{0}_{n}}|_{[o,p_{k_{i}}]})>0$.
Proof of Claim 1. For a contradiction, assume that this is not the case. Hence, there exists $k_{i}$ such that $\ell(\alpha_{o,w^{0}_{n}}|_{[o,p_{k_{i+1}}]})-\ell(\alpha_{o,w^{0}_{n}}|_{[o,p_{k_{i}}]})\leq 0$. If this is the case, then
Subclaim 1.
There exists $j\geq k_{i+1}$ such that $\ell(\alpha_{o,w^{0}_{n}}|_{[p_{j},p_{k_{i}}]})\leq 10C_{1}^{2}D\log(n)$.
Proof of 1 Let $q>k_{i+1}$ be the minimal index such that $\ell(\alpha_{o,w^{0}_{n}}|_{[o,p_{q}]})-\ell(\alpha_{o,w^{0}_{n}}|_{[o,p_{k_{i}}])}>0$ (which exists as $k_{i+1}<n$). If $\ell(\alpha_{o,w^{0}_{n}}|_{[p_{q},p_{k_{i}}]})\leq 10C_{1}^{2}\log(n)$ then $j=q$ proves the subclaim. If $\ell(\alpha_{o,w^{0}_{n}}|_{[p_{q-1},p_{k_{i}}]})\leq 10C_{1}^{2}\log(n)$ then $j=q-1$ proves the subclaim. If neither holds, then by $K$-bounded jumps of $(w^{*}_{n})_{n}$ we get
$K$ $\geq d(w^{o}_{q-1},w^{o}_{q})$
$\geq D^{-1}\ell(\alpha_{o,w^{0}_{n}}|_{[p_{q-1},p_{q}]})-D-d(p_{q-1},w^{o}_{q-1})-d(p_{q},w^{o}_{q})$
$\geq 20C_{1}^{2}\log(n)-D-2C_{1}^{2}\log(n),$
a contradiction. This finishes the proof of 1$\diamondsuit$ We are assuming that the event $\mathcal{B}$ holds, hence as $|j-k_{i}|\geq|k_{i+1}-k_{i}|\geq C_{3}\log(n)$ we must have that $d_{S}(w^{o}_{j},w^{o}_{k_{i}})>|j-k_{i}|/C_{3}\geq 100C_{1}^{2}D^{2}\log(n)$. On the other hand, by Subclaim 1, we also have that $d_{S}(w^{o}_{j},w^{o}_{k_{i}})\leq\ell(\gamma_{j})+\ell(\gamma_{k_{i}})+d_{S}(p_{j},p_{k_{i}})\leq 2C_{1}\log(n)+10C_{1}^{2}D^{2}\log(n)+D$, by the fact that $\alpha$ is a $D$-quasi-geodesic. A contradiction, thus proving Claim 1. $\blacksquare$
Hence, the lengths $(\ell(\alpha_{o,w^{0}_{n}}|_{[o,p_{k_{i}}]}))_{k_{i}}$ are increasing. We further note that if $\mathcal{B}$ holds, then $d_{S}(w^{o}_{k_{i}},w^{o}_{k_{i+1}})\geq 100D^{2}C_{1}^{2}\log(n)$ and so $d_{S}(p_{k_{i}},p_{k_{i+1}})\geq 98C_{1}^{2}\log(n)$. Thus under the assumption that events $\mathcal{A},\mathcal{B}$ both hold we get, by Claim 1 and $K$-bounded jumps of the Markov chain:
$\begin{split}Kn&\geq d_{S}(o,w^{o}_{n})\geq\frac{1}{D}\ell(\alpha)-1\\ &\geq\frac{1}{D}\sum_{0\leq i\leq t}(\ell(\alpha_{o,w^{0}_{n}}|_{[p_{k_{i}},p_{k_{i+1}}]}))-1\\ &\geq\frac{t}{D}\Big{(}\frac{98C_{1}^{2}\log(n)}{D}-D\Big{)}-1.\end{split}$
Hence, $t\leq\frac{D^{2}(Kn+1)}{98C_{1}^{2}\log(n)-D}$. Further, for all $k_{i}$, consider loops $P_{k_{i}}$ to be defined by the concatenation of $\gamma^{-1}_{k_{i}}$, $[w_{k_{i}},w_{k_{i+1}}],\gamma_{k_{i+1}}$, and $\alpha_{o,w^{0}_{n}}|_{[p_{k_{i}},p_{k_{i+1}}]}$. The length of each loop $P_{k_{i}}$ is at most $300KC_{1}^{2}C_{3}D^{2}\log(n)$ and hence the area of each $P_{k_{i}}$ is at most $\delta_{G}(300KC_{1}C_{3}D^{2}\log(n))$. By filling all of them, we get a filling for the null-homotopic word $W_{o,w_{1}^{o},\ldots,w^{o}_{n}}$ (defined as in Definition 2.9).
Hence, in the case that the events $\mathcal{A}$ and $\mathcal{B}$ holds, we get that
$\begin{split}\mathbb{E}\Big{[}\text{Fill}_{\mathcal{P},\alpha}(o,\dots,w^{o}_{n})\Big{|}\mathcal{A}\cap\mathcal{B}\Big{]}&\leq\sum_{0\leq i\leq t}\operatorname{Area}_{\mathcal{P}}(P_{k_{i}})\\ &\leq t\delta_{G}(300KC_{1}^{2}C_{3}D^{2}\log(n))\\ &\leq\frac{D^{2}(Kn+1)}{98C_{1}^{2}\log(n)-D}\delta_{G}(300KC_{1}^{2}C_{3}D^{2}\log(n)).\end{split}$
And thus, by the law of total expectation
$\begin{split}\mathbb{E}\Big{[}&\text{Fill}_{\mathcal{P},\alpha}(o,\dots,w^{o}_{n})\Big{]}\\ &=\mathbb{E}\Big{[}\text{Fill}_{\mathcal{P},\alpha}(o,\dots,w^{o}_{n})\Big{|}\mathcal{A}\cap\mathcal{B}\Big{]}\mathbb{P}[\mathcal{A}\cap\mathcal{B}]+\mathbb{E}\Big{[}\text{Fill}_{\mathcal{P},\alpha}(o,\dots,w^{o}_{n})\Big{|}\mathcal{A}^{C}\cup\mathcal{B}^{C}\Big{]}\mathbb{P}[\mathcal{A}^{C}\cup\mathcal{B}^{C}]\\ &\leq\frac{D^{2}(Kn+1)}{98C_{1}^{2}\log(n)-D}\delta_{G}(300KC_{1}^{2}C_{3}D^{2}\log(n))+((K+D)n+D)^{d}(C_{1}n^{1-C_{1}}+n^{-d-1})\\ &\leq C\Big{(}\frac{n}{\log(n)}\delta_{G}(\log(n))\Big{)}+Cn+C\end{split}$
for some $C$ big enough, where we use the fact that for two events $A,B$ we have $\mathbb{P}\Big{[}A\cap B\Big{]}\geq 1-\mathbb{P}[A^{C}]-\mathbb{P}[B^{C}]$ and use Theorem 3.2, Lemma 3.1 and the choice of $C_{1}$ to conclude.
This bound is independent of $o$ and hence taking the supremum over all $o\in G$ concludes the proof of Theorem 3.3, by the usual order of functions defined above. ∎
As a consequence of Theorem 3.3 we obtain the following two consequences and C, which we state again for the reader’s convenience.
Corollary 3.4 (A).
Let $G$ be a finitely presented acylindrically hyperbolic group with Dehn function $\delta_{G}$ at most polynomial. Let $(Z_{n})_{n}$ be a simple random walk on $G$. For every quasi-geodesic combing $\alpha$ of $G$, we have that
$R\delta_{(G,(Z_{n}),\alpha)}(n)\preccurlyeq\frac{n}{\log(n)}\delta_{G}(\log n).$
In particular, $R\delta_{(G,(w^{p}_{n}),\alpha)}(n)\preccurlyeq\delta_{G}(n)$.
Proof.
This follows immediately from Theorem 3.2 and Theorem 3.3 by noting that the random walks considered are tame Markov chains by Lemma 2.5. ∎
Corollary 3.5 (B).
Let $G$ be one of the following:
  • A non-elementary hyperbolic group;
  • a non-elementary relatively hyperbolic group;
  • an acylindrically hyperbolic $3$-manifold group;
  • the fundamental group of a graph of groups with nilpotent undistorted edge groups and at most polynomial Dehn function
Let $(w_{n}^{\ast})$ be a tame Markov chain on $G$. Then, for any quasi-geodesic combing $\alpha$ of $G$ we have that:
$R\delta_{(G,(w^{*}_{n}),\alpha)}(n)\preccurlyeq\frac{n}{\log n}\delta_{G}(\log(n)).$
In particular, $R\delta_{(G,(w^{p}_{n}),\alpha)}(n)\prec\delta_{G}(n)$.
Proof.
This follows from Theorem 3.2 as all these groups satisfy either [GS22, Assumption 1 or Assumption 2] and hence we can apply Theorem 3.3 by noting that all acylindrically hyperbolic $3$-manifold groups have quadratic Dehn function (by the fact that they have an automatic structure [ECH+92, Theorem 12.4.7] as they do not contain $Sol$ or $Nil$ components in their geometric decomposition), to conclude. ∎
Corollary 3.6 (C).
Let $G$ be one of the following:
  • the mapping class group of a finite type surface;
  • an extra-large Artin group;
  • or more generally a well-behaved hierarchically hyperbolic group (in the sense of [GHP+23]).
Let $(w_{n}^{\ast})$ be a tame quasi-homogeneous Markov chain on $G$. Then, for any quasi-geodesic combing $\alpha$ of $G$ we have that:
$R\delta_{(G,(w^{*}_{n}),\alpha)}(n)\preccurlyeq n\log n.$
In particular, $R\delta_{(G,(w^{p}_{n}),\alpha)}(n)\prec\delta_{G}(n)$.
Proof.
We know that tame quasi-homogenenous Markov chains satisfy the conclusion of Theorem 3.2. Hence, applying Theorem 3.3 and using the fact fact that hierarchically hyperbolic groups have quadratic Dehn function [BHS19, Corollary 7.5] concludes the proof. ∎